# multi-sr.bib

### v1

@@ 5 Vandermonde matrices in the super-resolution regime
In this section we come back to the super-resolution problem, and consider the Vandermonde matrices $U(x_{1},\dots,x_{n};{\cal G}_{N})$ in (3) in the single cluster setting. We prove the multidimensional analogue of [batenkov2021, Theorem 2.3], with the caveat that the geometry of the nodes stays fixed as $\text{SRF}\gg 1$. We start by defining the cluster geometry. Naturally, since the sampling set is an integer grid, we must restrict the nodes to $\mathbb{T}^{d}:=[-\pi,\pi)^{d}\equiv\left(\mathbb{R}\mod 2\pi\right)^{d}$ to avoid aliasing.
Definition 10.
@@
Further, let ${\cal G}_{N}=\{0,\pm 1,...,\pm N\}^{d}$ be the uniform symmetric grid of side length $12N+$ 1.
@@
where $t_{k}:=\operatorname{rank}(V_{\leq k}({\cal X}))-\operatorname{rank}(V_{\leq k-1}({\cal X}))$, $k=1,\dots,m$.
@@
With the identification $D_{N}({\cal X})=K_{\epsilon}({\cal Y})$ as above, we would like to apply Theorem 1 and Corollary 1 to the kernel $\mathcal{K}=\mathcal{K}_{N}$. However, $\mathcal{K}$ depends on $N$, so we need to make sure that all the estimates in the proofs are uniform in $N$.
Definition 12 (sinc kernel).
@@
Outline of the proof: we aim to prove that the eigenvalues of $D_{N}$ converge to $\operatorname{Sinc}$ kernel when $N\to\infty$. The scaling of the eigenvalues of $\operatorname{Sinc}$ can be obtained using Corollary 1 (the constants $\tilde{\lambda}_{k,j}$ don’t depend on $N$). First we show element-wise convergence of $D_{N}$ to $\operatorname{sinc}$ and equicontinuity of $D_{N}$ in $\epsilon<\epsilon_{0}$ to show uniform convergence of $D_{N}$ to $\operatorname{sinc}$ (uniform in $\epsilon<\epsilon_{0}$ and independent of $N$). Then using the Bauer–Fike Theorem, we prove convergence of the eigenvalues.
@@
Theorem 6 (The Bauer–Fike Theorem).
@@
Definition 13 (Element-wise convergence).
@@
Definition 15 (Uniform convergence).
@@
The first direction is an immediate corollary of Theorem 7.25 in [rudin1964]. Since the sequence converges pointwise, we can replace the subsequence in the proof of (b) by the sequence itself. The second direction is Theorem 7.24. ∎
Lemma 5.
@@
We proved that each entry of $D_{N}({\cal X})$ converges point-wise to each entry of $\operatorname{Sinc}_{\epsilon}(\mathcal{Y})$. Now we show that this convergence is uniform for $\epsilon\in[0,\epsilon_{0}]$ for a given $\epsilon_{0}$. We show that using Theorem 7. First we need to show that the sequence $\{\big{(}D_{N}({\cal X})\big{)}_{i,j}\}$ is equicontinuous.
@@
Using Lemma 5 and 6 together with Theorem 7 we get the following Corollary.
Corollary 2.
@@
We take $\nu$ in Definition 14 to be the minimum of all the $\nu_{i,j}^{(\ell)}$ according to the proof of Lemma 6, with $\ell=1,\dots,d$. Note that for $y_{i,j}^{(\ell)}=0$ we have $\nu_{i,j}^{(\ell)}=\infty$. ∎
Corollary 3 (Uniform convergence of eigenvalues).
@@
Since $\operatorname{Sinc}_{\epsilon}(\mathcal{Y})$ is a symmetric matrix, the eigenvectors matrix $E$ is orthogonal in the eigenvalue decomposition of $\operatorname{Sinc}_{\epsilon}(\mathcal{Y})$ i.e $\operatorname{Sinc}_{\epsilon}(\mathcal{Y})=E\Lambda{E}^{-1}$. Thus we have $\kappa_{2}(E)=1$. Using the Bauer–Fike theorem 6 with Corollary 2 and $\delta:={\delta^{\prime}\over n\sqrt{n}}$, we get:
@@
To complete the proof of Theorem 5, we choose $\delta^{\prime}=O\biggl{(}(\alpha\epsilon_{0})^{2m}\biggr{)}$ in (17) and apply Corollary 1. The only remaining question is to show that $\det(RWR)\neq 0$. We will show that the limit $N\to\infty$ the rank of $W=\lim_{N\to\infty}\frac{1}{(2N)^{d}}W_{N}$ is full, and as a result, $\det\left(\tilde{R}W\tilde{R}^{T}\right)$ is uniformly bounded away from zero. By definition of the Wronskian, we write
@@
[…]
@@
[…]
@@
[…]
Since the set of monomials is linearly independent, we have that $W$ is positive definite. Thus $W$ is full rank, and as a result, $\det\left(\tilde{R}W\tilde{R}^{T}\right)>0$. This completes the proof of Theorem 5. ∎
Remark 6.
During the final stages of preparation of our manuscript, we became aware that Weilin Li is working on a related topic [li2024a]. Relative to our results, more explicit conditions and constants are derived for the scaling of the smallest singular value of the multivariate Vandermonde matrix (we consider the entire spectrum); multi-cluster geometry is investigated; only cube and spherical sampling sets are considered; the decay rate in the case of uniform dilations of generic sets is not sharp; finally, a form of geometric characterization condition is used to describe the node geometry (as opposed to the sequence $\{t_{k}\}$ in our work).
@@ 6.1 Dirichlet kernel eigenvalue scaling
We start with the numerical study of the eigenvalues of the Dirichlet matrix $D_{\epsilon}({\cal X},{\cal G}_{n})$, confirming the scaling predicted by Lemma 3. Let $d=2$, $n=6$ and $\mathcal{X}=\{(x_{i},y_{i})\}_{i=1}^{n}$. We checked two scenarios:
@@ 6.2.1 Local Stability
Analogous to some previous studies in the one-dimensional case, e.g. [batenkov2013c, batenkov2018-genspikes], we consider the “local stability” of the problem to be well-approximated by the error incurred by the nonlinear least squares (NLS) method applied to the noisy measurements:
@@
We have implemented the 2D ESPRIT algorithm according to [sahnoun2017, Section IV-A]. The experimental setup was as follows:
@@ Appendix A Taylor Expansion
Let $\mathcal{K}\in\mathcal{C}^{(r,r)}(\Omega)$ where $\Omega$ is an open neighborhood of 00 and $r\geq m+1$, where $m=\mu({\cal X})$ is the discrete moment order of ${\cal X}$ (Definition 5). We recall the Maclaurin expansion for bivariate functions from [usevich2021, section 5]. For $\epsilon x,\epsilon y$ such that $[0,\epsilon x],[0,\epsilon y]\subset\Omega$ (where $[0,x]$ is a line segment from 00 to $x$), we have:
@@ Appendix B Auxiliary Lemmas
Recall formula (0.8.12.3) for sum of matrices in [horn2013]:

### v2

@@ 5 Vandermonde matrices in the super-resolution regime
In this section we come back to the super-resolution problem, and consider the Vandermonde matrices $U(x_{1},\dots,x_{n};{\cal G}_{N})$ in (3) in the single cluster setting. We prove the multidimensional analogue of [undefg, Theorem 2.3], with the caveat that the geometry of the nodes stays fixed as $\text{SRF}\gg 1$. We start by defining the cluster geometry. Naturally, since the sampling set is an integer grid, we must restrict the nodes to $\mathbb{T}^{d}:=[-\pi,\pi)^{d}\equiv\left(\mathbb{R}\mod 2\pi\right)^{d}$ to avoid aliasing.
Definition 5.1.
@@
Further, let ${\cal G}_{N}=\{0,\pm 1,...,\pm N\}^{d}$ be the uniform symmetric grid of side length $2N+1$.
@@
where $t_{k}:=\operatorname{rank}(V_{\leq k}({\cal X}))-\operatorname{rank}(V_{\leq k-1}({\cal X}))$, $k=1,\dots,m$. Here $\asymp$ means ’exact scaling’ as defined in Corollary 3.5.
@@
With the identification $D_{N}({\cal X})=K_{\epsilon}({\cal Y})$ as above, we would like to apply Theorem 3.1 and Corollary 3.5 to the kernel $\mathcal{K}=\mathcal{K}_{N}$. However, $\mathcal{K}$ depends on $N$, so we need to make sure that all the estimates in the proofs are uniform in $N$.
Definition 5.4 (sinc kernel).
@@
Outline of the proof: we aim to prove that the eigenvalues of $D_{N}$ converge to $\operatorname{Sinc}$ kernel when $N\to\infty$. The scaling of the eigenvalues of $\operatorname{Sinc}$ can be obtained using Corollary 3.5 (the constants $\tilde{\lambda}_{k,j}$ don’t depend on $N$). First we show element-wise convergence of $D_{N}$ to $\operatorname{sinc}$ and equicontinuity of $D_{N}$ in $\epsilon<\epsilon_{0}$ to show uniform convergence of $D_{N}$ to $\operatorname{sinc}$ (uniform in $\epsilon<\epsilon_{0}$ and independent of $N$). Then using the Bauer–Fike Theorem, we prove convergence of the eigenvalues.
@@
Theorem 5.5 (The Bauer–Fike Theorem).
@@
Definition 5.6 (Element-wise convergence).
@@
Definition 5.8 (Uniform convergence).
@@
The first direction is an immediate corollary of [undefai, Theorem 7.25]. Since the sequence converges pointwise, we can replace the subsequence in the proof of (b) by the sequence itself. The second direction is Theorem 7.24. ∎
Lemma 5.10.
@@
We proved that each entry of $D_{N}({\cal X})$ converges point-wise to each entry of $\operatorname{Sinc}_{\epsilon}(\mathcal{Y})$. Now we show that this convergence is uniform for $\epsilon\in[0,\epsilon_{0}]$ for a given $\epsilon_{0}$. We show that using Theorem 5.9. First we need to show that the sequence $\{\big(D_{N}({\cal X})\big)_{i,j}\}$ is equicontinuous.
@@
Using Lemma 5.10 and 5.11 together with Theorem 5.9 we get the following Corollary.
Corollary 5.12.
@@
We take $\nu$ in Definition 5.7 to be the minimum of all the $\nu_{i,j}^{(\ell)}$ according to the proof of Lemma 5.11, with $\ell=1,\dots,d$. Note that for $y_{i,j}^{(\ell)}=0$ we have $\nu_{i,j}^{(\ell)}=\infty$. ∎
Corollary 5.13 (Uniform convergence of eigenvalues).
@@
Since $\operatorname{Sinc}_{\epsilon}(\mathcal{Y})$ is a symmetric matrix, the eigenvectors matrix $E$ is orthogonal in the eigenvalue decomposition of $\operatorname{Sinc}_{\epsilon}(\mathcal{Y})$ i.e $\operatorname{Sinc}_{\epsilon}(\mathcal{Y})=E\Lambda{E}^{-1}$. Thus we have $\kappa_{2}(E)=1$. Using the Bauer–Fike theorem 5.5 with Corollary 5.12 and $\delta:={\delta^{\prime}\over n\sqrt{n}}$, we get:
@@
To complete the proof of Theorem 5.3, we choose $\delta^{\prime}=O\biggl((\alpha\epsilon_{0})^{2m}\biggr)$ in (18) and apply Corollary 3.5. The only remaining question is to show that $\det(RWR)\neq 0$. We will show that the limit $N\to\infty$ the rank of $W=\lim_{N\to\infty}\frac{1}{(2N)^{d}}W_{N}$ is full, and as a result, $\det\left(\tilde{R}W\tilde{R}^{T}\right)$ is uniformly bounded away from zero. By definition of the Wronskian, we write
@@
$\lim_{N\to\infty}\prod_{j=1}^{d}\sum_{k=-N}^{N}\frac{1}{2N}\frac{(i)^{\alpha_{j}}(-i)^{\beta_{j}}}{\alpha_{j}!\beta_{j}!}\biggl({k\over N}\biggr)^{\alpha_{j}+\beta_{j}}={1\over 2^{d}}\int_{[-1,1]^{d}}{{(i)^{|\alpha|}(-i)^{|\beta|}}\over{\alpha!\beta!}}x^{\alpha+\beta}\,dx$
@@
$W=\lim_{N\to\infty}\frac{1}{(2N)^{d}}W_{N}=\bigg[{1\over 2^{d}}\int_{[-1,1]^{d}}{{(i)^{|\alpha|}(-i)^{|\beta|}}\over{\alpha!\beta!}}x^{\alpha+\beta}\,dx\bigg]_{\alpha,\beta\in\mathbb{P}_{m}}.$
@@
$W_{\alpha,\beta}=\langle x^{\alpha},x^{\beta}\rangle={1\over 2^{d}}\int_{[-1,1]^{d}}{{(i)^{|\alpha|}(-i)^{|\beta|}}\over{\alpha!\beta!}}x^{\alpha+\beta}\,dx.$
Since the set of monomials is linearly independent, we have that $W$ is positive definite. Thus $W$ is full rank, and as a result, $\det\left(\tilde{R}W\tilde{R}^{T}\right)>0$. This completes the proof of Theorem 5.3. ∎
Remark 5.14.
During the final stages of preparation of our manuscript, we became aware that Weilin Li is working on a related topic [undefx]. Relative to our results, more explicit conditions and constants are derived for the scaling of the smallest singular value of the multivariate Vandermonde matrix (we consider the entire spectrum); multi-cluster geometry is investigated; only cube and spherical sampling sets are considered; the decay rate in the case of uniform dilations of generic sets is not sharp; finally, a form of geometric characterization condition is used to describe the node geometry (as opposed to the sequence $\{t_{k}\}$ in our work).
@@ 6.1 Dirichlet kernel eigenvalue scaling
We start with the numerical study of the eigenvalues of the Dirichlet matrix $D_{\epsilon}({\cal X},{\cal G}_{n})$, confirming the scaling predicted by Lemma 4.5. Let $d=2$, $n=6$ and $\mathcal{X}=\{(x_{i},y_{i})\}_{i=1}^{n}$. We checked two scenarios:
@@ 6.2.1 Local Stability
Analogous to some previous studies in the one-dimensional case, e.g. [undefa, undefc], we consider the “local stability” of the problem to be well-approximated by the error incurred by the nonlinear least squares (NLS) method applied to the noisy measurements:
@@
We have implemented the 2D ESPRIT algorithm according to [undefaj, Section IV-A]. The experimental setup was as follows:
@@ Appendix A Taylor Expansion
Let $\mathcal{K}\in\mathcal{C}^{(r,r)}(\Omega)$ where $\Omega$ is an open neighborhood of 0 and $r\geq m+1$, where $m=\mu({\cal X})$ is the discrete moment order of $\cal X$} (Definition 2.5). We recall the Maclaurin expansion for bivariate functions from [undef, section 5]. For $\epsilon x,\epsilon y$ such that $0,\epsilon x],[0,\epsilon y]\subset\Omega$ (where $[0,x]$ is a line segment from 0 to $x$), we have:
@@ Appendix B Auxiliary Lemmas
Recall formula (0.8.12.3) for sum of matrices in [undefs]: